OpenStudy (anonymous):

Find the slope-intercept equation that passes through (-1, 2) and is perpendicular to the line with the equation 6y - x = 12. A. y = -1/6x + 11/6 B. y = 1/6x + 13/6 C. y = 6x + 8 D. y = -6x -4

OpenStudy (anonymous):

for this you first need to get your equation into y = mx + b form. do you know how to do that?

OpenStudy (anonymous):

so 6y = x + 12 now isolate the y by dividing both sides by 6.

OpenStudy (anonymous):

so the equation of the line they give you is y = 1/6x + 2 right?

OpenStudy (anonymous):

the perpendicular line has a slope that is m1 = -1/m2

OpenStudy (anonymous):

whats m1 and m2 ?

OpenStudy (anonymous):

m1 is the slope of the first line.. so -1/m2 is the slope of the second line.

OpenStudy (anonymous):

so the slope is the negative reciprocal of the original slope.

OpenStudy (anonymous):

so if your slope is 1/6 the slope of the perpendicular line would be -6 or -6/1 .. the negative reciprocal

OpenStudy (anonymous):

so now you have the slope of the line 2.. you can plug in the points they give you into the point slope formula y - y1 = m(x-x1) to get the equation of the line you are looking for

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

yes i think so . let me try and work it out!

OpenStudy (anonymous):

ok :) lemme know what u come up with.

OpenStudy (anonymous):

i got y = -4x -4 ?

OpenStudy (anonymous):

-6x since -6 is the slope.

OpenStudy (anonymous):

-6(x + 1) + 2 = -6x - 6 + 2 = -6x - 4

OpenStudy (anonymous):

oh i added 2 to the -6. oops :)

OpenStudy (anonymous):

great! good luck :) medal my answer if it helped

OpenStudy (anonymous):

i was planning on it! haha thank you (: